OpenStudy (lena772):

A reaction has a rate law of : rate = k [ClO3-] [Cl-] [H+]2. If the initial reaction rate is 3.2E-3 M/s, then what is the rate constant if the initial concentrations are [ClO3-]o= 0.10 M, [Cl-]o= 0.10 M, [H+]o= 0.20 M ? (units are M-3 " s-1) A) 1.6 B) 8.0 C) 1.2 D) 0.62 E) none of these

OpenStudy (lena772):

@Abhisar

OpenStudy (abhisar):

Ok, so have you tried it first urself?

OpenStudy (lena772):

i think you would say 3.2 E-3 M/s = k [0.10] [0.10] [0.2] k= (3.2E-3/([0.10]*[0.1)]*[0.20])

OpenStudy (abhisar):

Yes, that's correct c:

OpenStudy (lena772):

k= 0.064

OpenStudy (abhisar):

Check for one thung!!!

OpenStudy (lena772):

so i guess D iis closest

OpenStudy (abhisar):

No, you did a minor mistake !!!

OpenStudy (lena772):

actually wrong brackets... 1.6

OpenStudy (lena772):

A?

OpenStudy (abhisar):

\(\sf Rate = k[ClO_3^-][Cl^-][H^+]^2\)

OpenStudy (abhisar):

So, K should be equal to ???

OpenStudy (lena772):

1.6*0.10*0.20*0.10=3.2 E-3 (the initial rate) so k is 1.6

OpenStudy (lena772):

A :) thank you

OpenStudy (lena772):

That was incorrect :(

OpenStudy (abhisar):

Nuuuuu!!! \(\sf K=\Large \frac{Rate}{[ClO_3^-][Cl^-][H^+]^2}\)

OpenStudy (abhisar):

You need to keep the order as given, I said you did a mistake. look above.. :(

OpenStudy (lena772):

i got 1.6 when i calclulated that

OpenStudy (abhisar):

It should be 8

OpenStudy (abhisar):

\(\sf K=\Large \frac{3.2 \times 10^{-3}}{0.1\times 0.1 \times (0.2)^2}\)

OpenStudy (lena772):

why is 0.2 squared

OpenStudy (abhisar):

Because the Rate Law is given as \(\sf Rate = k[ClO_3^-][Cl^-][H^+]^2\)

OpenStudy (lena772):

oh okay I didn't see that !

OpenStudy (abhisar):

C;

OpenStudy (lena772):

thank you!